40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Apr 15 18:59:41 2011, tara, DailyProgress, FSS, Preparing to install TTFSS 
    Reply  Tue Apr 19 22:40:03 2011, tara, DailyProgress, FSS, Preparing to install TTFSS 
       Reply  Wed Apr 20 17:54:52 2011, tara, DailyProgress, FSS, Preparing to install TTFSS 000_0025.jpg000_0024.jpg000_0023.jpgfastout.png
          Reply  Sat Apr 23 11:40:09 2011, tara, DailyProgress, FSS, Preparing to install TTFSS TTFSS_cartoon.pdf
             Reply  Sun Apr 24 22:02:07 2011, tara, DailyProgress, FSS, Preparing to install TTFSS beat_2011_04_24.pngbeat_2011_04_24.png
                Reply  Mon Apr 25 17:53:20 2011, tara, DailyProgress, FSS, Preparing to install TTFSS beat_2011_04_25.png
Message ID: 589     Entry time: Sun Apr 24 22:02:07 2011     In reply to: 587     Reply to this: 591
Author: tara 
Type: DailyProgress 
Category: FSS 
Subject: Preparing to install TTFSS 

 I can lock the laser frequency to RCAV with TTFSS, but the noise in the beat is higher than before.

 I'm trying to bring the beat back. This is the to do list

1) check the Power from LO input to match the designed value for TTFSS (10dB)

   The LO power available is 20 dBm

  The mixer, JMS-1H, requires 10dBm input. I used 10 dB Attenuator to attenuate the LO signal before plugging in the servo.

   I forgot to use the attenuator before. I'm not sure if the 20dBm signal will affect the RF amplifier and the others components or not. I might have to check the box again**.

 

2) check the error signal, make sure that the phase is right.

     To make sure that the box is working properly, I checked the error signal, the details are here. Since we changed the setup the phase adjustment between LO and PD needs to be changed. It is set to 5.3659 V to make the error signal symmetric.

 

3) optimize the gain, between common gain and fast gain. To minimize the noise from MIXER MON

     I adjusted common gain and fast gain while monitored the noise spectrum from MIXER MON from dc - 10 MHz to make sure that the noise was minimized and there was no oscillation at any frequency.

     COMMON GAIN knob is set to 455. FAST GAIN knob is set to 800.

 

4) Then I checked the beat signal. It is an order of magnitude higher than before.

 beat_2011_04_24.png

The setup:

1mW power into both cavities

 RCAV gain: FAST 800, COMMON 455.

ACAV gain: 7

 

Attachment 1: beat_2011_04_24.png  49 kB  Uploaded Mon Apr 25 00:50:51 2011  | Hide | Hide all
beat_2011_04_24.png
ELOG V3.1.3-